40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 5 09:48:57 2010, Koji, Update, SUS, ITMs placed on the tables in the chambers 
    Reply  Mon Apr 5 11:39:41 2010, steve, Update, SUS, ITM-south installation 
Message ID: 2769     Entry time: Mon Apr 5 11:39:41 2010     In reply to: 2766
Author: steve 
Type: Update 
Category: SUS 
Subject: ITM-south installation 


Steve and Koji (Friday, Apr 02)


Installation of ITMs are going on. Two new ITMs were placed on the optical table in the vacuum chambers. ITM for the south arm was put at the right place in accordance to the CAD drawing. ITM for the east arm is still at a temporaly place.

Tower placement (10:30-11:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

ITM (South arm) (14:00-16:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

- Leveled the table approximately.

- Released the EQ stops

- Removed anchors for the OSEM cables as it was too short. The wire distribution will be changed later.

- Put the OSEMs. Adjust the insertion to the middle of the OSEM ranges.

- Clamped the EQ stops again

- Placed the tower to the right place according to the CAD drawing.

- Released the EQ stops again.

- Check the OSEM values. The LL sensor showed small value (~0.5). Needs to be adjusted.


ITM (South) damping adjustment

- Found the signs for the facing magnets are reversed.

- Otherwise it damps very well.


 The cabling on the seismic stack was rerouted so it could reach the south edge of the table: the cables  were removed from the viton padded clamps and repositioned this morning.

ITM-south tower's earthquake screw viton tips could be a little bit larger. They do not stay in their screw hole after a hard clamping action.

4-40 earthquake screws under the test mass:viton tips can fall out without action, the treads are cross threaded so the screws are wobbling


ELOG V3.1.3-